# Dept. of Systems Engineering, ETS Montreal, Canada

### v1

@@ 3.1 Prototype Learning Module
Describing part occurrence to a single feature vector enables us to exchange these features for the corresponding prototype in different modalities and creates the intermediate step. Apart from the semantic feature contrastive of prototypes, for pushing more on prototype to point different semantic body-parts in different locations of images, the visual contrastive is used as:
$\mathcal{L}_{\text{vc}}=\sum_{k=1}^{K}\sum_{q=k+1}^{K}\sum_{u\in U}[\|M^{k}_{m}-M^{q}_{m}\|]_{u},\vspace{-0.15cm}$ (9)
which makes the model extract them from diverse regions from the backbone features with the minimum intersection.
@@
(c) Part Discrimination: The information encoded by prototypes must describe the object in the foreground. That is, the mutual information between the joint probability $\{P^{k}\}_{k=1}^{K}$ and identity $Y$ should be maximized:
$MI(P^{1},\dots,P^{K};Y).\vspace{-.2cm}$ (10)
@@
In the supplementary material, we proofed that for maximizing Eq. (10), we can minimize cross-entropy loss between each part-prototypes features and identities. In fact, by doing this, the model makes the part prototype regions lie on the object and have enough information about the foreground. In other words, each part prototype should be able to recover the identity of the person in the images. So part ID-discriminative loss is defined as:
$\mathcal{L}_{\text{pid}}=\frac{-1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C_{y}}y_{c}\log(W^{k}_{c}(\mathbf{p}_{m}^{k})),\vspace{-0.25cm}$ (11)
@@
where $W^{k}$ is linear layer predicting the probability of identity $y_{c}$ from $\mathbf{p}_{m}^{k}$. To avoid the prototype diversity, we do not share parameters of $W^{k}$ and $W^{q}$.
3.2 Bidirectional Multi-step Learning
The ultimate goal of V-I person ReID is to match individuals by extracting the features that not only discriminate the person attributes but are also independent of V or I modality. To extract discriminative features, the Attentive Prototype Embedding (APE) module uses the extracted part prototypes to represent the final features by weighting the significance of the parts.
(a) Attentive Prototype Embedding Instead of directly concatenating learned prototypes as the final feature, we process all other prototypes in input images. The motivation of this design is that each prototype may discriminate the part-attributes shared for all individuals rather than ID attributes. The Attentive Prototype Embedding module (which is detailed in Fig. 3), $\mathcal{F}$, leverages relevant person information between prototypes by applying an attention mechanism to achieve person-aware final features. To aggregate the ID-discriminative features of prototypes, the APE module first uses a fully connected layer to score and emphasize important channels in their feature map. It weights each prototype feature by computing the similarities between them.
$\begin{gathered}\mathcal{F}(\mathbf{A}_{m})=\mathbf{W}_{\text{mlp}}(\mathbf{C}_{m}),\text{ where }\\ \mathbf{C}_{m}=\mathbf{W}_{v}(\mathbf{A}_{m})\otimes\mathbf{B}_{m}\text{, }\mathbf{B}_{m}=\sigma(\mathbf{W}_{q}(\mathbf{A}_{m})\otimes\mathbf{W}_{k}(\mathbf{A}_{m})),\vspace{-0.25cm}\end{gathered}$ (12)
and
$\mathbf{A}_{m}=[\mathbf{p}_{m}^{1},\mathbf{p}_{m}^{2},\dots,\mathbf{p}_{m}^{K}],\vspace{-0.2cm}$ (13)
Figure 3: Attentive prototype embedding (APE) architecture.
(b) Bidirectional Multi-Step Learning To deal with the significant shift between modalities in the feature space, the model is trained via multiple intermediate steps that gradually bridge this domain gap. Using intermediates with less domain shift, the model gradually learns to leverage cross-modality discriminating clues at each step[39, 5, 1, 57]. In the beginning, the discrepancies between modalities are small letting the model learn from the easier samples first, then converge later on more complex cases with larger shifts. One solution to achieve these intermediates is to start from one modality as the source to transform gradually to the other as the target. However, by doing this, the model may forget the learned knowledge of the source and be biased on the target.
To address this issue, modalities are transformed bidirectionally to intermediate domain be created at each step. By gradually transforming domains to each other from V and from I, the domain gap vanishes over multiple steps. Each intermediate auxiliary step provides discriminative information about the person across both modalities, enabling the training process to transform inputs from one to the other. By creating these intermediate domains, our model aims to neither lose nor duplicate information from the main domains. For gradual transformation in each step, we exchange the features of the same prototypes from both modalities to create a mixed and virtual representation space with less domain discrepancies. These prototype-mixed intermediate feature spaces are used alongside V and I prototypes to provide APE along with the ability to extract common features from the modalities. For step $t$, which $t\in\{1,\dots,T\}$ and $T$ is the number of intermediate steps, the intermediate features for the same person are generated using a mixing function $\mathcal{G}(.,.,.)$ as:
$\mathbf{A}^{(t)}_{m}=\mathcal{G}(\mathbf{A}_{m},\mathbf{A}_{\tilde{m}},t),$ (15)
where $r(.,.,.)$ is a weighted random selector:
where $\mathcal{U}(0,1)$ is a uniform random generator. For intermediate step ($t<T$) we have $\mathbf{f}^{(t)}_{m}=[\mathcal{F}(\mathbf{A}^{(t)}_{m});\mathbf{g}_{m}]$ and in the last step $\mathbf{f}^{(T)}_{m}=[\mathcal{F}(\mathbf{A}^{(T)}_{m});\mathbf{g}_{\tilde{m}}]$ is used. The module $\mathcal{F}$ learns to gradually reduce the discrepancies in representation space by applying metric learning objectives bidirectionally between $\mathbf{f}_{i}$ and $\mathbf{f}^{(t)}_{i}$ and between $\mathbf{f}_{v}$ and $\mathbf{f}^{(t)}_{v}$ as follows:
$\begin{split}\mathcal{L}_{\text{bcc}}&=\mathcal{L}_{\text{cc}}(\mathbf{f}_{v},\mathbf{f}^{(t)}_{v})+\mathcal{L}_{\text{cc}}(\mathbf{f}_{i},\mathbf{f}^{(t)}_{i})\\ \mathcal{L}_{\text{bce}}&=\mathcal{L}_{\text{ce}}(\mathbf{f}_{v})+\mathcal{L}_{\text{ce}}(\mathbf{f}^{(t)}_{v})+\mathcal{L}_{\text{ce}}(\mathbf{f}_{i})+\mathcal{L}_{\text{ce}}(\mathbf{f}^{(t)}_{i})\\ \mathcal{L}_{\text{reid}}&=\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{bcc}},\end{split}\vspace{-0.2cm}$ (18)
where $\mathcal{L}_{\text{cc}}$, $\mathcal{L}_{\text{ce}}$ are center cluster loss and cross-entropy loss [46].
Since $\mathbf{f}^{(t)}_{i}$ begins from $\mathbf{f}_{i}$ at first steps and approaches to $\mathbf{f}_{v}$ at ending steps, the model is gradually trained to extract more robust modality invariant features from images by seeing samples with low modality gap to the harder ones with higher gap[39].
3.3 Training and Inference
The deep backbone model is trained over multiple steps, where in each step, the model tries to represent the same feature space for two input domains, V/I and the intermediate ones, using the overall loss:
$\mathcal{L}=\mathcal{L}_{\text{reid}}+\lambda_{\text{f}}(\mathcal{L}_{\text{lfc}}+\mathcal{L}_{\text{hfc}})+\lambda_{\text{v}}\mathcal{L}_{\text{vc}}+\lambda_{\text{p}}\mathcal{L}_{\text{part}}+\lambda_{\text{i}}\mathcal{L}_{\text{pid}}\vspace{-0.1cm}$ (19)
After training, the tail and head backbones are combined for the inference. Then, it extracts the prototypes and global features as $\mathbf{A}_{i}$, $\mathbf{A}_{v}$, $\mathbf{g}_{i}$ and $\mathbf{g}_{v}$ from given input image $x_{i}$ and $x_{v}$. Using Eq. (14), the $\mathbf{f}_{i}$ and $\mathbf{f}_{v}$ are computed by the APE. The matching score for a query input image is computed by cosine similarly.
4 Results and Discussion
In this section, we compare the proposed BDMG with SOTA person ReID methods, including global, part-based, and intermediate approaches. We also conduct extensive ablation studies to demonstrate the effectiveness of the proposed bi-directional learning as well as the impact of the number of part-prototypes and intermediate steps during learning. The supplementary materials detail our experimental methodology protocol and implementation of baselines, and performance metrics, and SYSU-MM01 [42] and RegDB [32] datasets.
@@ 4 Results and Discussion
BMDG (ours)a - 75.43 97.42 72.86 83.53 98.02 83.77 93.92 98.11 89.18 94.08 97.0 88.67
BMDG (ours)b - 76.39 97.90 78.22 83.59 98.96 85.87 94.76 98.91 92.21 94.56 98.31 93.07
@@ 4.1 Comparison with State-of-Art Methods:
Table 1 compares the accuracy of BMDG with state-of-the-art V-I ReID approaches. Our experiments on SYSU-MM01 and RegDB datasets show that BMDG outperforms these methods in various situations. This is due to BMDG ability to capture more ID-related knowledge across modalities by learning discriminative part-prototypes and gradually reducing the modality-specific information in extracted final features. The prototype alignment module also creates gradual and bidirectional intermediate spaces without sacrificing discriminative ability, allowing to outperform the intermediate family of methods. In addition, our approach has the advantage that it can be easily integrated into different part-based V-I ReID models to improve the generality by training gradually. However, in comparison to [20] on SYSU-MM01 dataset, we have lower R1 accuracy (-2%) in the All Search settings, in Indoor settings (+2%) and RegDB (+10%) we have higher performance.
4.2 Ablation Study:
(a) Step size and number of part prototypes. We evaluate the best option for the number of parts and steps. Table 2(a) shows that using our BMDG approach can increase the R1 accuracy when growing the number of steps $T$ for a specific number of prototypes $K$. This table shows performance increases for multi-step for specific $K$, which lets the model learn from the lower gap sample and converge on harder cases.
Table 2: Accuracy (R1%) of BMDG using (a) our prototype mixing and (b) Mixup [56] setting for different numbers of part prototypes ($K$) and intermediate steps ($T$). The mAP is reported in the supplementary materials.
(a) Prototype exchanging
@@ 4.2 Ablation Study:
(b) Mixup [56]
$T$ Number of part prototypes ($K$)
3 4 5 6 7 10
@@
Another option for each step, instead of exchanging some parts, is to use the Mixup [56] strategy for all parts from two modalities with normalized weight: $\mathbf{A}^{(t)}_{m}=\alpha^{(t)}\mathbf{A}_{m}+(1-\alpha^{(t)})\mathbf{A}_{m^{\prime}}$, where $\alpha$ starts from 0 to 1 w.r.t. to step $t$ and number of steps. Results in Table 2(b) show that using multi-step increases performance marginally since it does not utilize the ability of prototypes to create meaningful intermediate steps. This is explained by the fact that prototypes contain some modality-specific information, which is crucial for transforming domains in intermediate steps and may be lost in the averaging process.
(b) Part-based baselines. The effectiveness of BMDG is shown when integrated into state-of-the-art part-based or prototype-based baseline models [52, 46, 12]. We executed the original code published by authors for each method in our servers, using the hyper-parameters and other configurations they provided with and without BMDG. Table 3 shows that BMDG improves performance for mAP and rank-1 accuracy for all methods by average of 2.02% and 1.54 % respectively when integrated into these baselines.
Method R1 (%) mAP (%)
DDAG [52] 53.62 52.71
DDAG with BMDG 55.36 (+1.74$\uparrow$) 54.05 (+1.34$\uparrow$)
MPANet [46] 66.24 62.89
MPANet with BMDG 68.74 (+2.50$\uparrow$) 64.25 (+1.36$\uparrow$)
SAAI [12] 71.87 68.16
SAAI with BMDG 73.69 (+1.82$\uparrow$) 70.08 (+1.92$\uparrow$)
(c) Loss functions. Contrastive objectives between parts ($\mathcal{L}_{\text{lfc}}$ and $\mathcal{L}_{\text{hfc}}$) are essential for the model to detect similar semantics between objects, thereby creating an intermediate step by swiping corresponding parts. Table 4 shows this necessity when used in conjunction with $\mathcal{L}_{\text{part}}$ – this loss provides appropriate supervision to create steps and increase performance at test time. As can be seen, using such regularization in the multi-step approach increases mAP by 3% and R1 by 4%. In Table 4, the effect of part separation loss $\mathcal{L}_{\text{part}}$ is shown. Results indicate that using this loss pushes the parts to be scattered on the body regions and gives the part features a more robust representation. Also, it enables the intermediate steps to generalize the feature space better by increasing the performance over the single-step with multi-step setting. The part identity loss $\mathcal{L}_{\text{pid}}$ allows the model to extract ID-related prototypes features that contain discriminative information about the object and generate an id-informative intermediate step. Without this objective, parts features are potent for focusing on background information. As shown in Table 4 (compare the two last rows), performance increases when using this objective.
Table 4: Impact on the accuracy of using different BMGD losses, using SYSU-MM01 data under single-shot for single-step ($T=0$) and multi-step ($T=K$) settings. The baseline $\mathcal{L}_{\text{reid}}$ loss is used in all cases.
@@
$\mathcal{L}_{\text{pid}}$ $\mathcal{L}_{\text{lfc}}$ $\mathcal{L}_{\text{hfc}}$ $\mathcal{L}_{\text{vc}}$ $\mathcal{L}_{\text{part}}$ $T=0$ $T=K$ $T=0$ $T=K$
@@
(d) Bidirectional domain generalization. To show the impact of the BMGD approach, we conduct experiments under the following settings, with (1) single-step, (2) one-directional multi-step (from V $\rightarrow$ I and I $\rightarrow$ V), and (3) bidirectional for creating the intermediate series. The results are shown in Table 4.2. Although using multiple intermediate steps from one modality to another improves the model’s accuracy due to letting the model leverage common discriminative by gradually learning, this transition may forget about the source modality. So, the bi-directional approach deals with forgetting challenges and has the best performance since it is not biased on a specific modality.
Table 5: Impact on the rank-1 and mAP accuracy using different settings for multi-step domain generalization. All numbers are percent. $V\rightarrow I$ means that the intermediates are created only from V, by replacing prototypes from I features.
@@
|c||cc|cc|cc|cc| Settings Number of part prototypes ($K$)

### v2

@@ 3.1 Prototype Learning Module
Describing part occurrence to a single feature vector enables us to exchange these features for the corresponding prototype in different modalities and creates the intermediate step. Apart from making part features independent of each other by feature contrastive losses, for pushing prototypes to focus on different semantic body parts in different image locations, the visual contrastive is used as:
$\vspace{-0.15cm}\mathcal{L}_{\text{vc}}=\sum_{k=1}^{K}\sum_{q=k+1}^{K}\sum_{u\in U}[\|M^{k}_{m}-M^{q}_{m}\|]_{u},$ (10)
to extract parts from diverse regions of backbone features with the minimum intersection. Also, for equivariant matching of regions, a random rigid transformation ($R$) is applied on the input images ($x_{m}$) to create the transformed images ($\tilde{x}_{m}$). Both are then processed by the model, and the transformation on the masking score ($\tilde{M}_{m}$) is inverted from the transformed images. The objective is to minimize the distance between the transformed and original maps:
$\vspace{-0.15cm}\mathcal{L}_{\text{eq}}=\sum_{k=1}^{K}\|M^{k}_{m}-R^{-1}(\tilde{M}^{k}_{m})\|.$ (11)
@@
(c) Part Discrimination. The information encoded by prototypes must describe the object in the foreground. That is, the mutual information between the joint probability $\{P^{k}\}_{k=1}^{K}$ and identity $Y$ should be maximized:
$\text{MI}(P^{1},\dots,P^{K};Y).\vspace{-.2cm}$ (12)
@@
In the supplementary material, we proofed that for maximizing Eq. (12), we can minimize cross-entropy loss between each prototypes features and identities. In fact, by doing this, the model makes the part prototype regions lie on the object and have enough information about the foreground. In other words, each part prototype should be able to recover the identity of the person in the images. So part ID-discriminative loss is defined as:
$\mathcal{L}_{\text{p}}=\frac{-1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C_{y}}y_{c}\log(W^{k}_{c}(\mathbf{p}_{m}^{k})),\vspace{-0.25cm}$ (13)
@@
where $W^{k}$ is linear layer predicting the probability of identity $y_{c}$ from $\mathbf{p}_{m}^{k}$. To ensure prototype diversity and discrimination, we do not share parameters of $W^{k}$ and $W^{q}$ and randomly drop out a proportion of them during training, preventing any single part from dominating.
@@
BMDG (ours)a - 76.15 97.42 73.21 83.53 98.02 84.92 93.92 98.11 89.18 94.08 97.0 88.67
BMDG (ours)b - 78.08 97.90 78.22 83.59 98.96 86.35 94.76 98.91 92.21 94.56 98.31 93.07
@@
3.2 Bidirectional Multi-step Learning
After identifying prototypes, multiple auxiliary intermediate domains are created, and features are extracted using the Attentive Prototype Embedding (APE) module.
(a) Attentive Prototype Embedding. Instead of directly concatenating learned prototypes as the final feature, we process all other prototypes in input images. The motivation of this design is that each prototype may discriminate the part-attributes shared for all individuals rather than ID attributes. The Attentive Prototype Embedding module (which is detailed in the supply. materials), $\mathcal{F}$, leverages relevant person information between prototypes by applying an attention mechanism to achieve person-aware final features. To aggregate the ID-discriminative features of prototypes, the APE module first uses a fully connected layer to score and emphasize important channels in their feature map. It weights each prototype feature by computing the similarities between them.
$\begin{gathered}\mathcal{F}(\mathbf{A}_{m})=\mathbf{W}_{\text{mlp}}(\mathbf{C}_{m}),\text{ where }\mathbf{C}_{m}=\mathbf{W}_{v}(\mathbf{A}_{m})\otimes\mathbf{B}_{m}\text{, }\\ \mathbf{B}_{m}=\sigma(\mathbf{W}_{q}(\mathbf{A}_{m})\otimes\mathbf{W}_{k}(\mathbf{A}_{m})),\vspace{-0.25cm}\end{gathered}$ (14)
and
$\mathbf{A}_{m}=[\mathbf{p}_{m}^{1};\mathbf{p}_{m}^{2};\dots;\mathbf{p}_{m}^{K}],\vspace{-0.1cm}$ (15)
(b) Bidirectional Multi-Step Learning. To deal with the significant shift between modalities in the feature space, the model is trained via multiple intermediate steps that gradually bridge this domain gap. Using intermediates with less domain shift, the model gradually learns to leverage cross-modality discriminating clues at each step[33, 4, 1, 51]. Initially, the discrepancies between modalities are small, letting the model learn from the easier samples first, then converge on more complex cases with larger shifts. One solution to achieve these intermediates is to start from one modality as the source and transform gradually to the other as the target. However, this makes the model forget the learned knowledge of the source and be biased on the target.
To address this issue, modalities are transformed bidirectionally to create an intermediate domain at each step. By gradually transforming domains to each other, the domain gap vanishes over multiple steps. Each intermediate auxiliary step provides discriminative information about the person across both modalities, enabling the training process to transform inputs from one to the other. By creating these intermediate domains, our model aims to neither lose nor duplicate information from the main domains. For gradual transformation in each step, we exchange the features of the same prototypes from both modalities to create a mixed and virtual representation space with less domain discrepancies. These prototype-mixed intermediate feature spaces are used alongside V and I prototypes to provide APE along with the ability to extract common features from the modalities. For step $t$, which $t\in\{1,\dots,T\}$ and $T$ is the number of intermediate steps, the intermediate features for the same person are generated using a mixing function $\mathcal{G}(.,.,.)$ as:
$\mathbf{A}^{(t)}_{m}=\mathcal{G}(\mathbf{A}_{m},\mathbf{A}_{\tilde{m}},t),$ (17)
where $r(.,.,.)$ is a weighted random selector:
where $\mathcal{U}(0,1)$ is a uniform random generator. For intermediate step ($t<T$), we have $\mathbf{f}^{(t)}_{m}=[\mathcal{F}(\mathbf{A}^{(t)}_{m});\mathbf{g}_{m}]$ and in the last step $\mathbf{f}^{(T)}_{m}=[\mathcal{F}(\mathbf{A}^{(T)}_{m});\mathbf{g}_{\tilde{m}}]$ is used. The module $\mathcal{F}$ learns to gradually reduce the discrepancies in representation space by applying metric learning objectives bidirectionally between $\mathbf{f}_{i}$ and $\mathbf{f}^{(t)}_{i}$ and between $\mathbf{f}_{v}$ and $\mathbf{f}^{(t)}_{v}$ as :
$\begin{split}\mathcal{L}_{\text{bcc}}&=\mathcal{L}_{\text{cc}}(\mathbf{f}_{v},\mathbf{f}^{(t)}_{v})+\mathcal{L}_{\text{cc}}(\mathbf{f}_{i},\mathbf{f}^{(t)}_{i})\\ \mathcal{L}_{\text{bce}}&=\mathcal{L}_{\text{ce}}(\mathbf{f}_{v})+\mathcal{L}_{\text{ce}}(\mathbf{f}^{(t)}_{v})+\mathcal{L}_{\text{ce}}(\mathbf{f}_{i})+\mathcal{L}_{\text{ce}}(\mathbf{f}^{(t)}_{i})\\ \mathcal{L}_{\text{re}}&=\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{bcc}},\end{split}\vspace{-0.2cm}$ (20)
that $\mathcal{L}_{\text{cc}}$, $\mathcal{L}_{\text{ce}}$ are center cluster and cross-entropy losses [40]. Since $\mathbf{f}^{(t)}_{i}$ begins from $\mathbf{f}_{i}$ at first steps and approaches to $\mathbf{f}_{v}$ at ending steps, the model is gradually trained to extract more robust modality invariant features by seeing samples with low modality gap to the harder ones with higher gap[33].
3.3 Training and Inference
In each step, the model tries to represent the same feature space for two input domains, V/I and the intermediate ones, using the overall loss expressed as:
$\mathcal{L}=\mathcal{L}_{\text{re}}+\lambda_{\text{f}}(\mathcal{L}_{\text{lc}}+\mathcal{L}_{\text{hc}})+\lambda_{\text{v}}\mathcal{L}_{\text{vc}}+\lambda_{\text{c}}\mathcal{L}_{\text{c}}+\lambda_{\text{i}}\mathcal{L}_{\text{p}}+\lambda_{\text{e}}\mathcal{L}_{\text{eq}}\vspace{-0.1cm}$ (21)
where $\lambda$ are hyper-parameters for weighting losses. During inference, the tail and head backbones are combined to extract the prototypes and global features as $\mathbf{A}_{i}$, $\mathbf{A}_{v}$, $\mathbf{g}_{i}$ and $\mathbf{g}_{v}$ from given input image $x_{i}$ and $x_{v}$. Using Eq. (16), the $\mathbf{f}_{i}$ and $\mathbf{f}_{v}$ are computed by the APE. The matching score for a query input image is computed by cosine similarly.
4 Results and Discussion
In this section, we compare the BMDG with SOTA V-I ReID methods, including global, part-based, and intermediate approaches. Extensive ablation studies are conducted to show the effectiveness of the proposed bi-directional learning, the impact of the number of parts-prototypes, and intermediate steps during learning. The supplementary materials provide detailed information on our experimental methodology, including the implementation of baseline models, performance measures, and descriptions of the SYSU-MM01 [36], RegDB [27], and LLCM [52] datasets.
Family LLCM
V $\rightarrow$ I I $\rightarrow$ V
Method Venue R1 mAP R1 mAP
DDAG[46] ECCV’20 40.3 48.4 48.0 52.3
CAJ[45] ICCV’21 56.5 59.8 48.8 56.6
RPIG[2] ECCVw’22 57.8 61.1 50.5 58.2
DEEN[53] CVPR’23 62.5 65.8 54.9 62.9
BMDG (ours) - 63.4 66.3 56.4 63.5
Table 2: Accuracy of the proposed BMDG and state-of-the-art methods on the Large LLCM Dataset. All numbers are percent.
@@ 4.1 Comparison with State-of-Art Methods:
Table 1 compares the accuracy of BMDG with state-of-the-art V-I ReID approaches. Our experiments on the SYSU-MM01 and RegDB datasets show that BMDG outperforms the SOTA methods in the majority of situations. Compared to [18], BMDG has lower R1 accuracy (-1%) in the All Search settings on the SYSU-MM01 dataset. However, it achieves higher performance in Indoor settings (+2%) and on the RegDB dataset (+10%). These results can be attributed to BMDG’s ability to effectively capture more ID-related knowledge across modalities. This is achieved by learning discriminative part-prototypes and gradually reducing modality-specific information in the extracted final features. Additionally, the prototype alignment module creates gradual and bidirectional intermediate spaces without sacrificing discriminative ability, allowing BMDG to surpass intermediate methods. Moreover, BMDG can be easily integrated into different part-based V-I ReID models, enhancing generality through gradual training. To show its effectiveness, we integrated our approach into state-of-the-art part-based or prototype-based baseline models [46, 40, 12]. We executed the original code published by authors using the hyper-parameters and other configurations they provided with and without BMDG. Table 3 shows that BMDG improves performance for mAP and rank-1 accuracy for all methods by an average of 2.02% and 1.54 %, respectively. Similar performance improvements are observed on the large and complex LLCM dataset (see Table 2). Since LLCM was introduced recently, few papers reported their results, so we executed other approaches with published code on the LLCM dataset.
Method R1 (%) mAP (%)
DDAG [46] 53.62 52.71
DDAG with BMDG 55.36 (+1.74$\uparrow$) 54.05 (+1.34$\uparrow$)
MPANet [40] 66.24 62.89
MPANet with BMDG 68.74 (+2.50$\uparrow$) 64.25 (+1.36$\uparrow$)
SAAI [12] 71.87 68.16
SAAI with BMDG 73.69 (+1.82$\uparrow$) 70.08 (+1.92$\uparrow$)
4.2 Ablation Studies:
(a) Step size and number of part prototypes. We evaluate the best option for the number of parts and steps. Table 4 (top) shows that using the BMDG approach increases the R1 accuracy when increasing the number of steps $T$ for a specific number of prototypes $K$. Also, it increases performance in multi-step with $K$ parts, letting the model learn from the lower gap sample and converge on harder cases.
@@ 4.2 Ablation Studies:
Entire mixup (IDM [10])
@@
Table 4: R1 accuracy of BMDG using (a) our prototype mixing and (b) IDM [10] for different numbers of part prototypes ($K$) and intermediate steps ($T$). mAP is reported in suppl. materials.
Another option for each step, instead of exchanging some parts, is to use the Mixup[50] or IDM [10] strategy for all parts from two modalities with normalized weight: $\mathbf{A}^{(t)}_{m}=\alpha^{(t)}\mathbf{A}_{m}+(1-\alpha^{(t)})\mathbf{A}_{m^{\prime}}$, where $\alpha$ starts from 0 to 1 w.r.t. step $t$ and number of steps. Results in Table 4 (bottom) show that using multi-step increases performance marginally since it does not utilize the ability of prototypes to create meaningful intermediate steps. This is explained by the fact that prototypes contain some modality-specific information, which is crucial for transforming domains in intermediate steps and may be lost in the averaging process.
(c) Loss functions. Contrastive objectives between parts ($\mathcal{L}_{\text{lc}}$ and $\mathcal{L}_{\text{hc}}$) are essential for the model to detect similar semantics between objects, thereby creating an intermediate step by swiping corresponding parts. Table 5 shows this necessity when used in conjunction with $\mathcal{L}_{\text{c}}$. Using such regularization in the multi-step approach increases mAP by 3% and R1 by 4%. Also, results of part separation loss, $\mathcal{L}_{\text{vc}}$, indicate that using this loss pushes the parts to be scattered on the body regions and gives the part features a more robust representation. Additionally, it enables the intermediate steps to generalize the feature space better by increasing the performance over the single-step with multi-step setting. The part identity loss $\mathcal{L}_{\text{p}}$ lets the model extract ID-related prototype features that contain discriminative information about the object and generate an id-informative intermediate step. Without this objective, parts features are likely to focus on background information. As shown in the two last rows, performance increases when using this objective.
@@
$\mathcal{L}_{\text{p}}$ $\mathcal{L}_{\text{lc}}$ $\mathcal{L}_{\text{hc}}$ $\mathcal{L}_{\text{vc}}$ $\mathcal{L}_{\text{c}}$ $\mathcal{L}_{\text{eq}}$ $T=0$ $T=K$ $T=0$ $T=K$
@@
Table 5: Impact on the accuracy of using different BMDG losses, using SYSU-MM01 for single-step ($T=0$) and multi-step ($T=K$) settings. The baseline $\mathcal{L}_{\text{re}}$ loss is used in all cases.
(d) Bidirectional domain generalization. To show the impact of the BMGD approach, we conduct experiments under the following settings: (1) single-step, (2) one-directional multi-step (from V $\rightarrow$ I and I $\rightarrow$ V), and (3) bidirectional for creating the intermediate series. The results are shown in Table 4.2. Although using multiple intermediate steps from one modality to another improves the model’s accuracy by allowing it to gradually learn and utilize common discriminative features, this process can lead to losing information about the source modality. The bi-directional approach addresses these issues and achieves the best performance, as it is not biased towards any specific modality.
@@
—c——cc—cc—cc—cc— Settings Number of part prototypes ($K$)
@@
Table 6: Impact on rank-1 and mAP accuracy using different settings for multi-step domain generalization. All numbers are percent. $V\rightarrow I$ means that the intermediates are created only from V, by replacing prototypes from I features.